# Quantum Circuit Switching with One-Way Repeaters in Star Networks

### v1

@@ A-C Limits when $k\rightarrow\infty$
To conclude this Appendix, we show (informally in some cases) that the mean sojourn time converges to a finite value when the number of forwarding stations, $k$, goes to infinity. For that, we only need to show that the mean waiting time and the mean service time also converge to a constant.
@@
    where we have used the following: ($i$) $S_{i}$ follows a binomial distribution with $m$ attempts and probability of success $p$, ($ii$) $m!/(m-x)!<m^{x}$, and ($iii$) $\lim_{m\rightarrow\infty}m^{x}q^{m}$ = 0, for $x\geq 0$ and $0<q\leq 1$. Combining the previous result with (A.16), we obtain $\mathbb{E}[B_{n,w,p,m}]\rightarrow 1$ and $\mathbb{E}[B_{n,w,p,m}^{2}]\rightarrow 1$ when $k\rightarrow\infty$. Since $\mathbb{E}[T_{\mathrm{service}}]$ and $\mathbb{E}[T_{\mathrm{service}}^{2}]$ are linear in $\mathbb{E}[B_{n,w,p,m}]$ and $\mathbb{E}[B_{n,w,p,m}^{2}]$, they also converge to constant values.
@@ Appendix B Squared coefficient of variation of the service time
As discussed in Appendix A-A, there is no known general solution for the expected waiting time of an M/G/$s$ queue, with $s>1$. Approximations such as (A.11) work well when the squared coefficient of variation of the service time, $C_{\mathrm{service}}^{2}$, takes small values – in [33], the authors consider large values of $C_{\mathrm{service}}^{2}$ are in the order of $\gtrsim 10-100$. In this Appendix, we empirically show that $C_{\mathrm{service}}^{2}$ takes small values in systems in which the service time is linear in the number of batches until success of a window problem (this is the case for our QCS protocols).
@@
where $x$ and $y$ are non-negative constants and $B_{n,w,p,m}$ is the number of batches until success, as defined in (A.16). In our work, $x=2L/c$ is the travel time of a quantum data packet from user to user and $y=t_{\mathrm{fwd}}(2N+1)$ is the total forwarding time that a packet experiences from user to user. The first and second moments of the service time are
@@
Figure 7: The squared coefficient of variation of the window problem is generally below 1. Squared coefficient of variation of $B_{n,w,p,m}$, $C_{n,w,p,m}^{2}\equiv\mathrm{Var}\left[B_{n,w,p,m}\right]/\mathbb{E}\left[B_{n,w,p,m}\right]^{2}$. In the parameter regions explored, $C_{n,w,p,m}^{2}$ decreases with increasing window size $w$ and with increasing probability of success $p$. The behavior with $n$ and $m$ is nontrivial although the values remain below 1 (in the parameter regions explored). Each data point is the average over 10 simulations (error bars correspond to the standard error). Each simulation consisted in the estimation of $C_{n,w,p,m}^{2}$ over $10^{3}$ samples of $B_{n,w,p,m}$ collected via Monte Carlo sampling.
@@ Appendix C Parameter values
For the number of users $u$, the number of repeaters $N$, and the number of forwarding stations per repeater $k$, we employ values that are reasonable for early quantum networks, on the order of magnitude of $1-10$. Additionally, the specific values we choose provide illustrative examples (e.g., if the number of users was increased beyond 20, in general no service is possible for the other parameter values chosen). The physical size of the network, $L$, is in the order of magnitude of $1-10$ km. A short distance (1 km) is required for the forwarding stations based on the proposal from ref. [11], while all-photonic repeaters are expected to work at internode distances of $1-10$ km [22, 31].
The forwarding time of each station $t_{\mathrm{fwd}}$ is 100 $\mu$s in all our examples. We obtain this value if we consider that: ($i$) approximately ten gates (order of magnitude) are needed to forward a quantum data packet (e.g., the encoding and decoding circuits of the five-qubit perfect code consist of at most 8 gates each [36]), and ($ii$) a quantum gate takes around 1-50 $\mu$s (this is the case for qubits realized with color centers, such as nitrogen vacancies – see, e.g., refs. [37, 38]). The probability of success $p$ depends on $L$, on $N$, and also on the hardware used to implement the forwarding stations, as discussed in the main text. We assume the physical channels between users are optical fibers in which the speed of light is approximately $0.2$ km/$\mu$s. As discussed in the main text, we consider negligible control time since this would not affect our analysis.
Regarding the requests, increasing $n$ would increase the expected service time. We tested different values of $n$ and did not find any different behavior from the system. We chose a relatively small value of $n$ for our examples, as a large $n$ would increase the runtime of the Monte Carlo sampling when the analytical solutions cannot be applied. The request window $w$ must be as large as $n$ (otherwise sequential distribution is not possible and we can only distribute packets in parallel). We consider values of $w$ that are close to $n$ (7, 8, and 10) and also $w\rightarrow\infty$. We do not consider large values of $w$ since the results converge quickly to the infinite window case due to large values of $p$ being used (in most of our examples, $p\geq 0.7$). Lastly, the value of $\lambda_{0}$ was chosen to illustrate interesting behavior, for the other parameter values chosen. If $\lambda_{0}$ is much smaller, the system can process requests seamlessly and saturates at a larger number of users (i.e., the critical number of users increases). If $\lambda_{0}$ is much larger, the opposite happens. As a reference, each pair of users submits a request every $10^{4}$ $\mu$s (on average): during this time, a single forwarding station can forward 100 quantum data packets.
@@ Appendix E Critical number of users with probabilistic packet delivery ($p<1$)
In this Appendix, we provide more examples of the behavior of the critical number of users, $u_{\mathrm{crit}}$ with increasing number of forwarding stations, $k$, in the large-budget use case from Section III. Figure 9 shows $u_{\mathrm{crit}}$ vs $k$ for $w=7,8,10,\infty$ and $N=0,1,5$ (corresponding to $L=7.5,13,30$ km). In these cases, the probability of successful packet delivery is $p\approx 0.7$. As in the small-budget scenario, $u_{\mathrm{crit}}$ increases with increasing $k$ for sequential distribution of packets. Conversely, $u_{\mathrm{crit}}$ reaches a maximum value when packets are distributed in parallel, i.e., we cannot increase $u_{\mathrm{crit}}$ indefinitely by increasing $k$, for parallel distribution.
As discussed in the main text, parallel distribution supports more users than sequential distribution only when the window size is small (i.e., close to the number of states requested $n$) and when there are few forwarding stations per repeater (small $k$). In Figure 9, we have $n=7$, and we observe this behavior when $w=7,8$ but not for $w\geq 10$.
@@ Appendix F Mean sojourn time with all-photonic repeaters
Here, we provide some more examples to support the claims from Section III-B.
First, Figure 10 shows the mean sojourn time (MST) for sequential and parallel distribution of packets, for increasing number of users. In Figure 10a ($p=1$, $N=0$, $L=1$ km, as in Figure 4a, and $k=7$), parallel distribution always provides a lower MST but sequential distribution can support more users. In Figure 10b ($p\approx 0.7$, $N=0$, $L=7.5$ km, as in Figure 4b, and $k=2$), parallel distribution can support more users (up to 6, while the MST with sequential distribution diverges after 5 users).
Second, we provide the difference in MST between sequential and parallel distribution vs the number of users and the number of forwarding stations, for the large-budget use case from the main text. In this use case, we consider the all-photonic forwarding stations from ref. [31]. The number of intermediate repeaters between each user and the central repeater, $N$, is chosen to minimize the cost $N/(Lp)$. As discussed in the main text, the optimal solution is $p\approx 0.7$: for $L=7.5,13,30$ km, we have $N=0,1,5$. In Figure 11, we provide the results for window sizes $w=7,8,10,\infty$. The main features from this Figure are discussed in the main text.
Figure 10: Parallel distribution of packets is generally faster, and sometimes it supports more users. MST with sequential (blue) and parallel (orange) packet distribution, for increasing number of users $u$. (a) Small-budget use case ($p=1$; $N=0$, $L=1$ km) with $k=7$ and (b) large-budget use case ($p\approx 0.7$) with $N=0$, $L=7.5$ km, and $w=8$, and with $k=2$. Parameters used in this figure: $n=7$, $\lambda_{0}=10$ $\mu$s-4, $t_{\mathrm{fwd}}=100$ $\mu$s. MST in (a) calculated with (7). MST in (b) calculated with Monte Carlo sampling with $10^{5}$ samples (the error bars show the standard error).
@@
Appendix G Many users over long distances: more examples
In this Appendix, we provide more examples for the discussion in Section III-C. We provide evidence that all the discussion about the relation between critical distance, number of users, and number of repeaters holds in the parameter regions explored and not only in the example shown in the main text. In particular, Figures 12 and 13 show that the critical distance decreases with increasing number of users, for both parallel and sequential distribution of data packets, respectively. Moreover, increasing the number of repeaters $N$ does not always provide larger critical distances.
Note that the window size $w$ and the number of forwarding stations $k$ have a stronger effect on $L_{\mathrm{crit}}$ when packets are distributed sequentially: the differences between subplots are more noticeable in Figure 12 (sequential distribution) than in Figure 13 (parallel distribution).
Figure 12: Networks with many users cannot cover long distances when packets are distributed sequentially. Critical distance, $L_{\mathrm{crit}}$, vs number of users, $u$, for different numbers of repeaters $N$ in a star network. Users request $n=7$ quantum data packets to be successfully delivered within a time window $w=\infty,10,8,7$ (top to bottom). There are $k=6,12,18$ forwarding stations per repeater (left to right). Other parameters used in this figure: $\lambda_{0}=10$ $\mu$s-4, $c=0.2$ km/$\mu$s, $t_{\mathrm{fwd}}=100$ $\mu$s, $t_{\mathrm{control}}=0$. Results calculated using (8). For $w\rightarrow\infty$, $\mathbb{E}\left[B_{n,w,p,m}\right]$ (required to compute (8)) was computed using the analytical solution from Appendix A-B – in the other cases, the probability distribution of $B_{n,w,p,m}$ was estimated with $10^{4}$ Monte Carlo samples.
Figure 13: Networks with many users cannot cover long distances when packets are distributed in parallel. Critical distance, $L_{\mathrm{crit}}$, vs number of users, $u$, for different numbers of repeaters $N$ in a star network. Users request $n=7$ quantum data packets to be successfully delivered within a time window $w=\infty,10,8,7$ (top to bottom). There are $k=6,12,18$ forwarding stations per repeater (left to right). Other parameters used in this figure: $\lambda_{0}=10$ $\mu$s-4, $c=0.2$ km/$\mu$s, $t_{\mathrm{fwd}}=100$ $\mu$s, $t_{\mathrm{control}}=0$. Results calculated using (8). For $w\rightarrow\infty$, $\mathbb{E}\left[B_{n,w,p,m}\right]$ (required to compute (8)) was computed using the analytical solution from Appendix A-B – in the other cases, the probability distribution of $B_{n,w,p,m}$ was estimated with $10^{4}$ Monte Carlo samples.

### v2

@@ A-C Limits when $k\rightarrow\infty$
To conclude this Appendix, we show that the mean sojourn time converges to a finite value when the number of forwarding stations, $k$, goes to infinity. For that, we only need to show that the mean waiting time and the mean service time also converge to a constant.
@@
    where we have used the following: ($i$) $S_{i}$ follows a binomial distribution with $m$ attempts and probability of success $p$, ($ii$) $m!/(m-x)!<m^{x}$, and ($iii$) $\lim_{m\rightarrow\infty}m^{x}q^{m}$ = 0, for $x\geq 0$ and $0<q\leq 1$. Combining the previous result with (A.16), we obtain $\mathbb{E}[B_{n,w,p,m}]\rightarrow 1$ and $\mathbb{E}[B_{n,w,p,m}^{2}]\rightarrow 1$ when $k\rightarrow\infty$. Since $\mathbb{E}[T_{\mathrm{service}}]$ and $\mathbb{E}[T_{\mathrm{service}}^{2}]$ are linear in $\mathbb{E}[B_{n,w,p,m}]$ and $\mathbb{E}[B_{n,w,p,m}^{2}]$ (see (A.14) and (A.15)), they also converge to constant values.
@@ Appendix B Squared coefficient of variation of the service time
As discussed in Appendix A-A, there is no known general solution for the expected waiting time of an M/G/$s$ queue, with $s>1$. Approximations such as (A.11) work well when the squared coefficient of variation of the service time, $C_{\mathrm{service}}^{2}$, takes small values – in [33], the authors consider large values of $C_{\mathrm{service}}^{2}$ to be in the order of $\gtrsim 10-100$. In this Appendix, we empirically show that $C_{\mathrm{service}}^{2}$ takes small values in systems in which the service time is linear in the number of batches until success of a window problem (this is the case for our QCS protocols).
@@
where $x$ and $y$ are non-negative constants and $B_{n,w,p,m}$ is the number of batches until success, as defined in (A.16). In our work, $x=2L/c+2t_{\mathrm{fwd}}N$, which accounts for the travel time of a quantum data packet from user to user and for the processing delays introduced by the repeaters, and $y=t_{\mathrm{fwd}}$, which accounts for the delays in between (batches of) packets. The first and second moments of the service time are
@@
Figure 7: The squared coefficient of variation of the window problem is below 1 for the relevant parameter regimes. Squared coefficient of variation of $B_{n,w,p,m}$, $C_{n,w,p,m}^{2}\equiv\mathrm{Var}\left[B_{n,w,p,m}\right]/\mathbb{E}\left[B_{n,w,p,m}\right]^{2}$. In the parameter regions explored, $C_{n,w,p,m}^{2}$ decreases with increasing window size $w$ and with increasing probability of success $p$. The behavior with $n$ and $m$ is nontrivial although the values remain below 1 (in the parameter regions explored). Each data point is the average over 10 simulations (error bars correspond to the standard error). Each simulation consisted in the estimation of $C_{n,w,p,m}^{2}$ over $10^{3}$ samples of $B_{n,w,p,m}$ collected via Monte Carlo sampling.Parameters used in each subfigure: (a) $n=7$, $m=3$; (b) $n=7$, $p=0.7$; (c) $p=0.7$, $m=3$; (d) $m=1$.
@@ Appendix C Parameter values
For the number of users $u$, the number of repeaters $N$, and the number of forwarding stations per repeater $k$, we employ values that are reasonable for early quantum networks, on the order of magnitude of $1-10$. Additionally, the specific values we choose provide illustrative examples in which the system exhibits interesting and/or useful behavior (e.g., if we used the same parameter values but the number of users was much larger than 20, the mean sojourn times would diverge to infinity and no service would be possible). The physical size of the network, $L$, is in the order of magnitude of $1-10$ km. A short distance (1 km) is required for the forwarding stations based on the proposal from ref. [11], while all-photonic repeaters are expected to work at internode distances of $1-10$ km [22, 31].
The forwarding time of each station $t_{\mathrm{fwd}}$ is 100 $\mu$s in all our examples. We obtain this value if we consider that: ($i$) approximately ten gates (order of magnitude) are needed to forward a quantum data packet (e.g., the encoding and decoding circuits of the five-qubit perfect code consist of at most 8 gates each [36]), and ($ii$) a quantum gate takes around 1-50 $\mu$s (this is the case for qubits realized with color centers, such as nitrogen vacancies – see, e.g., refs. [37, 38]). The probability of success $p$ depends on $L$, on $N$, and also on the hardware used to implement the forwarding stations, as discussed in the main text. We assume the physical channels between users are optical fibers in which the speed of light is approximately $0.2$ km/$\mu$[39]. As discussed in the main text, we consider negligible control time since this would not affect our analysis.
Regarding the requests, increasing the number of requested states, $n$, would increase the expected service time. We tested different values of $n$ and did not find any different behavior from the system. In our examples, we use $n=7$. We chose this relatively small value of $n$ because a large $n$ would increase the runtime of the Monte Carlo sampling when the analytical solutions cannot be applied. The request window $w$ must be as large as $n$ (otherwise sequential distribution is not possible and we can only distribute packets in parallel). We consider values of $w$ that are close to $n$ (7, 8, and 10) and also $w\rightarrow\infty$. We do not consider large values of $w$ since the results converge quickly to the infinite window case due to large values of $p$ being used (in most of our examples, $p\geq 0.7$). Lastly, the value of $\lambda_{0}=10^{-4}$ $\mu$s-1 used in our examples was chosen to illustrate interesting behavior, for the other parameter values chosen. If $\lambda_{0}$ is small, the system can process requests seamlessly and saturates at a larger number of users (i.e., the critical number of users increases). If $\lambda_{0}$ is large, the opposite happens. Note that, when $\lambda_{0}=10^{-4}$ $\mu$s-1, each pair of users submits a request every $10^{4}$ $\mu$s (on average): during this time, a single forwarding station can forward 100 quantum data packets.
@@ Appendix E Critical number of users with probabilistic packet delivery ($p<1$)
In this Appendix, we provide more examples of the behavior of the critical number of users, $u_{\mathrm{crit}}$, with increasing number of forwarding stations, $k$, in the large-budget use case from Section III. Figure 9 shows $u_{\mathrm{crit}}$ vs $k$ for $w=7,8,10,\infty$ and $N=0,1,5$ (corresponding to $L=7.5,13,30$ km). In these cases, the probability of successful packet delivery is $p\approx 0.7$. As in the small-budget scenario, $u_{\mathrm{crit}}$ increases with increasing $k$ for sequential distribution of packets. Conversely, $u_{\mathrm{crit}}$ reaches a maximum value when packets are distributed in parallel, i.e., we cannot increase $u_{\mathrm{crit}}$ indefinitely by increasing $k$, for parallel distribution.
As discussed in the main text, parallel distribution supports more users than sequential distribution only when the window size is small (i.e., close to the number of states requested, $n$) and when there are few forwarding stations per repeater (small $k$). In Figure 9, we have $n=7$, and we observe this behavior when $w=7,8$ but not for $w\geq 10$. We also observe that, in the parameter regimes explored, this behavior vanishes when increasing the size of the network: for $L=7.5$ km, parallel distribution can support more users than sequential distribution for $w=7,8$ and small $k$; but for $L=30$ km, parallel distribution cannot support more users for any values of $w$ and $k$.
@@ Appendix F Mean sojourn time with all-photonic repeaters
Here, we provide some additional examples of the behavior of the mean sojourn time (MST) in terms of the number of users, $u$, and the number of forwarding stations per repeater, $k$.
Figure 10 shows the MST for sequential and parallel distribution of packets, for increasing number of users and fixed $k$. In Figure 10a ($p=1$, $N=0$, $L=1$ km, as in Figure 4a, and $k=7$), sequential distribution can support one more user than parallel distribution (up to 14). Nevertheless, parallel distribution provides a lower MST, although the advantage decreases as the number of users increases, since we approach the divergence. In Figure 10b ($p\approx 0.7$, $N=0$, $L=7.5$ km, as in Figure 4b, and $k=2$), parallel distribution can support more users (up to 6, while the MST with sequential distribution diverges after 5 users). In this case, the advantage in MST provided by parallel distribution actually increases with increasing number of users, since the divergence happens earlier when packets are distributed sequentially.
In Figure 11, we provide the difference in MST between sequential and parallel distribution vs the number of users and the number of forwarding stations, for the large-budget use case from the main text. In this use case, we consider the all-photonic forwarding stations from ref. [31]. The number of intermediate repeaters between each user and the central repeater, $N$, is chosen to minimize the cost $N/(Lp)$. As discussed in the main text, the optimal solution is $p\approx 0.7$: for $L=7.5,13,30$ km, we have $N=0,1,5$. In Figure 11, we provide the results for window sizes $w=7,8,10,\infty$. As discussed in the main text and in Appendix E, parallel distribution can support more users than sequential distribution only when the resources are scarce, i.e., when $w$ is small compared to $n$ and $k$ is small, and when the network is small. In the figure, this effect happens for $w=7,8$ and $L=7.5,13$ km.
Figure 10: Parallel distribution of packets is generally faster, and sometimes it supports more users. MST with sequential (blue) and parallel (orange) packet distribution, for increasing number of users $u$. (a) Small-budget use case ($p=1$; $N=0$, $L=1$ km) with $k=7$ and (b) large-budget use case ($p\approx 0.7$) with $N=0$, $L=7.5$ km, and $w=8$, and with $k=2$. Parameters used in this figure: $n=7$, $\lambda_{0}=10$ $\mu$s-4, $t_{\mathrm{fwd}}=100$ $\mu$s. MST in (a) calculated with (7). MST in (b) calculated using a discrete-event simulation and Monte Carlo sampling with $10^{5}$ samples (the error bars show the standard error).
@@
Appendix G Many users over long distances: sequential vs parallel distribution
In this Appendix, we compare sequential vs parallel distribution of packets in terms of the critical distance. As discussed in Section III-C, increasing the number of repeaters only increases the critical distance when the number of users is small. Figure 5 shows an example of this phenomenon when using sequential distribution of packets (the same example is shown in Figure 12a for convenience). In Figure 12b, we show that the same conclusions are observed when packets are distributed in parallel. Interestingly, when the number of users is fixed, sequential distribution (Figure 12a) allows for larger distances between them (i.e., it provides larger $L_{\mathrm{crit}}$) than parallel distribution (Figure 12b). We also tested different combinations of parameters ($w=7,8,10$ and $k=6,12,18$) and observed very similar qualitative and quantitative behavior.
Figure 12: Networks with sequential distribution of packets can cover longer distances than with parallel distribution. Critical distance, $L_{\mathrm{crit}}$, vs number of users, $u$, for different numbers of repeaters $N$ in a star network. (a) Sequential distribution of packets; (b) parallel distribution of packets. Parameters used in this figure: $n=7$, $w\rightarrow\infty$, $k=12$, $\lambda_{0}=10^{-4}$ $\mu$s-1, $c=0.2$ km/$\mu$s, $t_{\mathrm{fwd}}=100$ $\mu$s. Results calculated using (8), where $\mathbb{E}\left[B_{n,w,p,m}\right]$ (required to solve (8)) was computed using the analytical solution from Appendix A-B.